### v1

# A Planar Huygens Antenna Utilizing Crossed Electric and Magnetic Dipoles

Abstract
The natural source of a magnetic dipole in antennas is typically an electrically small loop, which can be utilized in conjunction with an electric dipole to realize an electrically small Huygens’ antenna. However, these antennas suffer from low radiation efficiency and their theoretical directivity limit is 4.8 dBi. Magnetic dipoles with an electrical size are highly desirable for high-gain applications. This paper builds on the development of a magnetic dipole source that utilizes a slot positioned near a printed dipole with a length twice that of the slot. Such a combination of electric and magnetic dipoles yields a highly directive radiation pattern, resulting in a higher gain than a uniformly illuminated antenna of similar size. The prototype is designed to operate at 4.5 GHz, with a directivity of up to 8.37 dBi. The analytical, numerical, and measured results agree fully. This high-gain superdirective antenna is highly desirable due to its excellent features, including being low-profile, PCB compatible, and having a low-complexity feeding topology, compared to the existing approach of the two-element end-fire array.
Index Terms:
Compact, electric dipole, Huygens’ source, magnetic dipole, superdirective.I Introduction
The pursuit of superdirective antennas has been an enduring challenge for nearly a century. Harrington investigated the fundamental relationship between antenna size and gain using spherical wave functions [1]. The theoretical maximum gain for an antenna enclosed within a sphere of radius , known as Harrington’s limit, is expressed as
Here, the assumption is considered to relate to the physical size of the antenna, indicating the approximate transition of rapid and slowly varying waves of the Hankel function with being the wavenumber and the radius of the sphere in which the antenna can be enclosed. For large , the gain limit simplifies to , which corresponds to the gain of a uniformly illuminated circular aperture. Although Harrington’s limit is often associated with superdirectivity, a widely accepted definition by Hansen [2] states that a superdirective antenna exhibits greater directivity than a similar-sized antenna with uniform excitation. The directivity of a uniformly illuminated aperture is given by
where is the free-space wavelength, and A represents the physical aperture size.
Oseen et al. [3] first considered the theoretical possibility of superdirectivity from arbitrarily small sources, followed by Riblet [4]. Uzkov [5] demonstrated that a linear array of isotropic radiators could achieve a maximum directivity of . The is an extraordinary directivity that can be achieved compared to the directivity of ”N”, realized by isotropic radiators half a wavelength apart. Several attempts have been made to demonstrate a superdirective antenna array. Bloch et al. experimentally demonstrated four half-wavelength dipoles spaced 0.2 apart to form a superdirective array [6]. Altshuler et al. [7] demonstrated an increase in gain of 6 dB compared to a single element of two monopoles 0.2 apart. Other subsequent works include an electrically small end-fire array by O’Donnell et al. [8] and Yaghjian et al. [9], the 2-element superdirective element by Best et al. [10], and the 4-element superdirective array by Clemente et al. [11]. These demonstrations utilize end-fire configurations to place antenna elements in the array. This approach is susceptible to narrow tolerances and requires a delicate balance of phase and amplitude to achieve a superdirective radiation pattern.
The size restrictions stipulated by Harrington in (1) do not apply to electrically small antennas (ESAs), and it is theoretically possible to achieve infinite gain, depending on the excitation of several modes. However, the Harrington limit cannot be applied for ESAs because the assumption is not valid for ESAs. Gayi [12] studied the Q of the antennas and defined the maximum gain limit for the ESAs. Kwon [13] obtained the expression for electrically small crossed electric and magnetic dipoles and showed that their gain could be as high as 3 (4.8 dB). Yaghjian [14] discussed the maximum achievable gain for the Huygens’ array and the possible realization of Huygens’ radiators. Pigeon et al. [15] investigated the feasibility of maximizing the directivity of miniaturized directive antennas using available techniques.
A heuristic limit on antenna maximum directivity was recently developed by Kildal et al. [16] for which directivity was estimated by accounting for the available spherical modes on the surface of a minimum sphere. For the limiting case of electrically small antennas, the limit becomes 4.8 dB, equivalent to a small Huygens antenna. Later, Pigeon et al. [17] renormalized Harrington’s limit for ESAs by defining the effective radius , where is the radius of the sphere that can enclose the ESA. This is especially in line with Wheeler’s proposal for the effective sphere of radius around an ESA, which is known as radiansphere [18].
The development of Huygens’ sources is challenging due to the unavailability of magnetic dipoles and the need for impedance matching to both electric and magnetic dipoles simultaneously. Several practical prototypes were developed, such as those of Ziolkowski et al. [19, 20, 21, 22, 23], which used combinations of loop and wire, and another by Niemi et al. [24], which used dual omega particles. The common element in these Huygen antennas is the utilization of a loop as a magnetic dipole that could be excited when the circumference of the loop is around [25]. The availability of magnetic dipoles in electrically small loops makes them an attractive choice for developing electrically small Huygens’ antennas.
These electrically small Huygens’ antennas suffer from a low radiation efficiency. They are more suitable for low-frequency applications, where tolerances are high for small elements, such as those of the order of . The maximum directivity of these electrically small Huygens’ antennas is also limited to 4.8 dB and cannot be employed for high-gain applications, especially in the mmWave range, where higher path losses due to weather effects can severely distort the signals. Recently, Ziolkowski [26] numerically investigated the feasibility of a mixed multipole antenna at 30 GHz, utilizing a layered structure that achieves high gain. Later, a practical design was demonstrated as an electrically small Huygens’ quadruple antenna, exploiting two pairs of Huygens’ sources [27].
This paper proposes a novel Huygens’ antenna topology for high-gain radiation that can be easily scaled to any frequency. The contribution of this paper is two-fold. First, we demonstrated magnetic dipole-type radiation with a slot when it was placed closely next to the printed dipole of twice the slot size, operating in full-wave dipole mode. Secondly, a low-profile, PCB-compatible, highly directive Huygens’ antenna is presented by the in-phase action of a magnetic dipole and a full-wave electric dipole, approaching the Harrington maximum limit. Contrary to similar recent superdirective antennas such as end-fire antenna array [28] and hybrid end-fire antenna (Huygens quadruple) [27], the proposed Huygens’ antenna does not require any complex feeding, can operate with nearly 100 radiation efficiency, and can be easily scaled to higher gains using low permittivity dielectrics.
II Antenna’s Design
The Geometry of the proposed Huygens’ antenna is shown in Fig. 1. The antenna consists of a cylindrical disk made of Rogers TMM3 board with dielectric constant and loss tangent . The thickness of the disk is mm, having 0.5 oz copper layers on both sides. A slot of length mm and width of mm is etched from the top side. On the bottom side, two arms of the dipole are printed along the y-axis, separated by a gap mm for feeding. The length and width of each arm are mm and mm, respectively. The lumped port is used to numerically feed the dipole arm and is then replaced by an SMA connector in the fabricated prototype. To support the magnetic dipole mode of the slot, each arm of the printed dipole is nearly the same size as the slot, resulting in a printed dipole that is twice the size of the slot, with a total effective electrical length of 1. This leads to the formation of a full-wavelength electric dipole mode. The excitation of the electric dipole mode in the printed dipole induces the magnetic dipole mode in the cross slot. The two dipoles radiate in phase, resulting in very high directivity along the +z-axis and cancellation of radiation in the backward direction, achieving a high front-to-back ratio of nearly 12 dB.
The operating principle of a Huygens’ antenna can be understood from the interference of electric and magnetic fields, as shown in Fig. 2. The -directed electric fields for electric and magnetic dipoles are oriented in the same direction along the + axis and opposite along the - axis for broadside radiation along the + direction. Similarly, the -directed magnetic fields constructively interfere along the +-axis and destructively interfere along the --axis. This set-up of crossed electric and magnetic dipoles results in a directive pattern along the +-axis known as the Huygens’ antenna. It should be noted that if the polarity of the electric or magnetic dipoles is reversed, a directional pattern along can also be obtained with the same set of dipoles.
In light of the discussed theory, we further analyzed the electric and magnetic fields of the proposed antenna topology, as shown in Figs. 3(a) and (b) for the electric dipole and in Figs. 3(c) and (d) for the magnetic dipole. The electric field over the two arms of the printed dipole indicates the accumulation of positive charges on one arm and negative charges on the other, forming a typical pattern of an electric dipole around which the magnetic field is revolving. The higher strength of the electric and magnetic fields along + indicates the constructive interference of both dipoles to form Huygens’ source. The slot acts as a magnetic dipole, as confirmed by the formation of south and north poles along the -axis and the rotation of the electric field around the slot, which is the typical behavior of a magnetic dipole. The -directed magnetic dipole and the -directed electric dipole form Huygens’ antenna.
The current distribution is plotted at different phases(time instants) as shown in Fig. 4. The peak current intensities along and across the slot appear simultaneously, indicating that the two dipoles are in phase. The slot typically does not radiate like a magnetic dipole. However, in the current scenario, when the printed dipole is twice the size of the 0.5 slot, antiparallel currents are induced on the dipole arms and the metallic plane, providing conducive conditions to excite the magnetic dipole in the slot, as can be seen in Fig. 5. The induced magnetic dipole in the slot, combined with the printed electric dipole, forms Huygens’ source, giving rise to the directive pattern along + and nullifying the backward radiation. The magnetic dipole in the slot is artificially engineered, contrary to its natural existence in nearly loops, which makes it an attractive choice by boosting the gain not only because of Huygens’ mode, but also because of size-dependent gain. This feature is not available in electrically small Huygens’ sources.
III Results Discussion
The proposed Huygens’ antenna is numerically verified using high-frequency simulation software (HFSS) and then fabricated from Rogers TMM3 boards using a PCB prototyping machine. The SMA connector feeds the EM energy through a vector network analyzer (VNA) with the center pin of the SMA connected to one arm and the ground connected to the other arm. The fabricated and assembled boards are shown in Fig. 6 along with their simulated and measured reflection response. The antenna’s operating frequency is 4.472 GHz in the experiments and 4.505 GHz in the measurements. A deviation of 33 MHz is observed compared to simulations, mainly attributed to the slight additional length introduced by the center pin of the connector. The dips, particularly on the left side of the central frequency, are mainly caused by a balanced-to-unbalanced system, which can typically be resolved using a balun. The reflection coefficient below -50 dB indicates that the antenna is very well matched at the operating frequency with almost 100% energy coupled to the antenna.
Since a low dielectric of TMM3 is used for high radiation efficiency, a simulated radiation efficiency of 97% is obtained with a directivity of 8.37 dBi. Since there is no loss in impedance matching and minimal ohmic loss, the realized gain is close to the maximally achievable directivity at 8.27 dBi, only 0.1 dB lower than the maximum directivity. The antennas are experimentally characterized in a compact range anechoic chamber. The experimental setup is shown in Fig. 7. The antenna under test, i.e. Huygens’ antenna, is placed on a rotating column, which is then illuminated by the plane wave from the reflector. A standard horn antenna feeds the reflector. The simulated and measured 2D radiation pattern in the XZ and YZ planes is shown in Fig. 8. The experimental and numerical results are in good agreement. The peak gain obtained is nearly 1.6 dB less than the measured value, primarily due to the slight variation in the perfect excitation of electric and magnetic modes resulting from the additional length introduced by the connector pin. Minor inaccuracies can also be caused by the cable on the back of the antenna, affecting the backward cancellation of the fields. Furthermore, the reference gain antenna has a tolerance of 0.5 dB.
The 3D plot for directivity is shown in Fig. 9. The peak directivity reaches 8.37 dBi, accompanied by a front-to-back ratio of approximately 12 dB. The high front-to-back ratio indicates near-equal matching with electric and magnetic dipoles. The overall antenna size of 1 increases the gain to a much higher value than could be achieved by the Huygens’ ESA. Due to the symmetrical nature of the fields in the front and back directions for Huygens’ source, the broadside direction can be easily switched by reversing the polarity of either the electric or magnetic fields. In the current design, backward radiation occurs at 3.86 GHz, a frequency slightly different from the operating frequency of 4.45 GHz. The maximum realized directivity of 8.37 dBi is 120% of that of the uniformly illuminated antenna of the same size calculated using (2). The maximum directivity with such a size can be 10.11 dBi, according to Harrington (1). The realized directivity is only 1.74 dB, less than the maximum Harrington limit.
IV Design Flexibility
The proposed approach can be easily scaled to any frequency as long as the ratio of 0.5 is maintained between the length of the slot and the total dipole length. Moreover, lower dielectric constants can enhance antenna gain as a result of increased electrical size. In this regard, three different dielectrics, that is, TMM3 with , TMM6 with , and RT/Duroid with are used as substrate to investigate the directivity in the context of the Harrington limit, keeping the overall physical size of the antenna the same, that is, it can fit in a sphere of radius 25 mm. The results for TMM3 are already presented in the previous section. The directivities are plotted in Fig. 10 for the other two. The antenna can achieve 9.33 dBi of directivity when RT/Duroid is used as a substrate because of the increased electrical size resulting from the frequency shift upward caused by the low permittivity of the dielectric. The maximum directivity corresponding to = 2.98, according to Harrington (1), is 11.7 dBi, introducing a gap of 2.37 dB between the realized and theoretical limits.
For the TMM6-based design, the frequency was shifted downward due to the high permittivity, decreasing the overall electrical size. The achieved directivity is approximately 7.2 dBi and the maximum directivity with this antenna, corresponding to = 1.849 according to Harrington (1), is 8.52 dBi, introducing a gap of 1.32 dBi between the achieved and Harrington limits. For the TMM3-based design, the maximum directivity in the forward mode is 10.118 dBi, corresponding to = 2.358. The achieved value is 8.37 dBi, resulting in a difference of approximately 1.74 dB. For the TMM3 backward mode, the maximum directivity is 9.07 dBi for , and the realized directivity is 7.4 dBi, resulting in a difference of approximately 1.67 dB. From here, it can be safely concluded that directivity approaches the Harrington limit as approaches 1. For 1, the updated limits should be used, as discussed in the introduction. For 1, the directivity approaches that of the uniformly illuminated antenna, as given by (2) and also discussed by Harrington [1].
All of the prototypes demonstrated so far are based on SMA mounted directly on the dipole. This could lead to backflow currents over the connector to the coaxial line, which is typically addressed using a balun. To avoid the balun, we proposed an alternative topology using edge-mounted SMA. The fabricated prototype and their return loss and directivity are shown in Fig. 11. A precise match of frequency response can be observed between the numerical and experimental results, except for a slight frequency shift due to minor fabrication tolerances. It should be noted that edge matching does not affect the field distribution, and thus essentially operates in Huygens’ mode, as described in detail earlier. Moreover, a good agreement is observed between the simulated and experimental gains, as depicted in 11. The measured gain peak is merely 0.25 dB lower than the numerically evaluated peak gain. The difference can be attributed to tolerance of the reference antenna, inaccuracy of alignment, and interference from the feeding cable with the pattern.
Furthermore, the proposed Huygens’ antenna is compared to available superdirective antennas in the literature as presented in Table I. The most popular approach to achieving superdirectivity is an end-fire array with closely spaced elements fed with specially designed signals, which requires a complex setup consisting of attenuators, amplifiers, phase shifters, and complex impedance matching networks. Although this approach demonstrates higher directivity, it still suffers from low radiation efficiency, resulting in a lower realized gain. So far, the two-element end-fire arrays presented in [28, 29] have demonstrated high radiation efficiency. The other approach involves simultaneously exciting multiple modes to achieve superdirectivity. Examples of which are the Huygens’ quadrupole antenna [27] and the unidirectional mixed multipole antenna, as numerically evaluated in [26].
| Topology | a | D | Complexityb | |||||
| (GHz) | (dBi) | (dBi) | () | (dBi) | ||||
| [7] | 2-element monopole array | 0.4 | 7.11 | 18.1 | 9.5 | - | - | High |
| [11] | 4-element dipole array | 0.868 | 1.7584 | 6.61 | 11.7 | - | -18 | High |
| [30] | 2-element magnetic dipole array | 0.435 | 3.28 | 12.35 | 9.9 | 55 | - | High |
| [29] | 3- and 4-element dipole array | 0.85, 0.85 | 1.46, 2.06 | 5.077, 8.36 | 9, 10.3 | - | -1.6, 2.1 | High |
| [31] | 2-element printed half loop array | 0.901 | 1.22 | 6 | 6.8 | 40 | - | High |
| [28] | 2-element strip dipole array | 3.5 | 1.55 | 7.44 | 6.5 | 99.3 | 6.3 | moderate |
| [26] | UMMA (Simulation only) | 28.3 | 2.11 | 9.38 | 9.52 | 77.9 | - | moderate |
| [27] | Huygens’ Quadrupole ESA | 1.662 | 0.98 | 4.65 | 7.61 | 88.1 | 7.06 | Low |
| Back Fed | Huygens’ Antenna | 4.5 | 2.3578 | 10.1 | 8.37 | 97 | 8.27 | Low |
| (This work) | ||||||||
| Edge Fed | Huygens’ Antenna | 4.1025 | 2.15 | 9.5 | 7.45 | 94 | 7.18 | Low |
| (This work) |
a The Harrington limit is used as the maximum gain; for an End-Fire array, another limit is based on the number of elements.
b The use of extra-impedance matching networks and attenuators/amplifiers is assumed to constitute high complexity; only phase shifters/multi-layer designs form moderate complexity, while simple fabrication and no extra circuits form low complexity.
It should be noted that the ESAs are not compared except in [27] because the Harrington limit does not apply to these. For Huygens’ ESAs, the maximum directivity limit is 4.8 dBi, which can only be surpassed by the excitation of multipoles [27, 26]. In all these approaches, the gain cannot be scaled to a higher gain than 7.06 dBi for the Huygens’ quadrupole antenna [27], except when bulky structures or large arrays are introduced. Here, using a low dielectric constant substrate such as RT/Duroid, we demonstrated a gain of 9.33 dBi, which could be even higher without using any substrate and can easily reach beyond 10 dBi.
V Conclusion
This paper proposes a superdirective antenna that utilizes crossed electric and magnetic dipoles, forming a Huygens antenna. A magnetic dipole is induced in the slot when it is placed near a printed dipole that is twice the size of the slot. The induced magnetic dipole interacts with the electric dipole mode of the printed dipole; destructive interference occurs in the backward direction, and constructive interference occurs in the forward direction, resulting in a highly directive pattern. The proposed Huygens’ antenna is designed to operate at 4.5 GHz, achieving a directivity of 8.37 dBi with a front-to-back ratio of 12 dB. The radiation efficiency of the antenna is 97%, corresponding to a gain achieved of 8.27 dBi. The high realized gain achieved from a compact planar structure with a high front-to-back ratio makes it suitable for several high-gain applications.